# Disentangling representations of retinal images with generative models

### v2

@@ 4.1 Learning disentangled subspaces with an encoder model
Our experiments showed that the age and camera subspaces were heavily entangled in the baseline setting when disentanglement was not specifically enforced (Fig. 5a), such that camera information was present in the age subspace and vice versa. Entanglement was visible by relatively high classification accuracy across subspaces (Fig.5a, off-diagonals). For example, camera type could be decoded from the age subspace $w_{\text{age}}$ with a performance of $28\%$ above chance level, compared to $45\%$ in the camera subspace. When enforcing disentanglement by additionally minimizing the distance correlation between subspaces the cross-subspace classification performance (off-diagonals in Fig. 5b) dropped, while the within-subspace classification was preserved (diagonals). In addition, the distance correlation between subspaces on the test data dropped as a direct consequence of our disentanglement loss (Fig. 5c). We could also observe the disentanglement effect qualitatively in 2D t-SNE visualizations of the learned representations. In both the baseline and the disentangled model, the age subspace exhibited structure w.r.t. age, and the camera subspace encoded camera type well (diagonals in Fig. 5d, e). In the baseline model, however, the age subspace also strongly encoded camera information (first row, second column in d).
Figure 7: Encoder disentanglement performance for ethnicity-camera setup. We trained encoder models on retinal images, where the first subspace $w_{\text{ethn}}$ encoded ethnicity $y_{\text{ethn}}$, and the second subspace $w_{\text{cam}}$ encoded camera information $y_{\text{cam}}$. In panels a and b, we show accuracy improvement over chance level accuracy. As a baseline model in a, we only trained linear classifiers on top of subspaces. In b, we further disentangle the subspaces using our distance correlation loss. Panel c compares the distance correlation measure between subspaces of our two model configurations. In panels d and e, we provide t-SNE visualizations comparing a baseline model with a model with disentanglement loss, respectively.
@@ 4.2 Ablation study for disentanglement loss
In our last ablation experiment, we explored the loss term $\lambda_{DC}$ for distance correlation minimization (Fig. 8c1-c3). As expected, the distance correlation measure between subspaces dropped with larger $\lambda_{DC}$ (Fig. 8c1). At the same time, the decoding accuracy for age from the age subspace (Fig. 8c2, blue curve) collapsed for $\lambda_{DC}\geq 2$, leading to a drop in decoding performance below 10%. In contrast, camera decoding (Fig. 8c2, green curve) from the age subspace becomes less accurate as soon as the disentanglement loss is introduced, indicating that the proposed method does what it is supposed to. For the camera subspace in (Fig. 8c3), the camera decoding accuracy remains high (green), and the age decoding accuracy initially drops (blue), in line with the goals of the distance correlation loss. However, the ablation study showed that the hyperparameter optimization of $\lambda_{DC}$ is essential, as over-weighting can lead to a complete collapse of the subspace encoding (Fig. 8c2,c3).
@@ 4.3 Learning disentangled subspaces with a generative model
Figure 9: Disentanglement performance of a generative model for age-camera setup. We trained generative models on retinal images, featuring three subspaces: $w_{\text{age}}$ encoding the age class $y_{\text{age}}$, $w_{\text{cam}}$ encoding the camera class $y_{\text{cam}}$, and $w_{\text{id}}$ serving as a patient-identity subspace. In a and b, we report kNN classifier accuracy improvement over chance level accuracy. The baseline model in a involved training linear classifiers on the first two subspaces. In b, we additionally disentangled the subspaces using our distance correlation loss. Panel c compares the distance correlation measure between subspaces of our two model configurations. Panels d and e present t-SNE visualizations of one baseline model and one trained generative model with the disentanglement loss, respectively.
@@
As with the predictive model, as a baseline method, we learned subspaces by appending linear classifiers, addressing distinct classification tasks in each subspace. In the generative framework, we introduced an additional latent subspace dedicated to encoding the patient’s identity $w_{\text{id}}\in\mathbb{R}^{16}$ (Fig. 4), enhancing our ability to generate high-resolution and realistic retinal images. To ensure independence between these subspaces, we incorporated the minimization of distance correlation. Consequently, we aimed to minimize the distance correlation among three distinct subspace combinations. Notably, we specifically enforced independence for the patient-identity space, ensuring that it remained uncorrelated with the subspaces encoding age, ethnicity, or camera information.
@@
Our disentanglement loss also showed an effect when applied to a generative image model (Fig. 9). While the baseline model for the generative model already demonstrated relatively good disentanglement performance, the introduction of the distance correlation loss further enhanced the separation of subspaces. For example, for the baseline model, age information could be decoded from the camera subspace and vice versa (Fig. 9a, off-diagonals). Additionally, age, and particularly camera information, were still contained within the identity subspace of the baseline model (last row, Fig. 9a). Upon incorporating the distance correlation loss, cross-subspace classification performance (off-diagonals and last row, Fig. 9b) dropped, while within-subspace classification was preserved (diagonals). Consequently, the distance correlation between subspaces on the test data decreased (Fig. 9c), underscoring the effect of our disentanglement loss. The qualitative impact of disentanglement was further evident in t-SNE visualizations: for example, the camera labels clustered in the patient identity space for the baseline model, but the clustering became more mixed up when we incorporated the distance correlation loss (Fig. 9d,e).
@@
Figure 10: Disentanglement performance of a generative model for ethnicity-camera setup. We trained generative models on retinal images, featuring three subspaces: $w_{\text{ethn}}$ encoding the ethnicity class $y_{\text{ethn}}$, $w_{\text{cam}}$ encoding the camera class $y_{\text{cam}}$, and $w_{\text{id}}$ serving as a patient-identity subspace. In a and b, we report kNN classifier accuracy improvement over chance level accuracy. The baseline model in a involved training linear classifiers on the first two subspaces. In b, we additionally disentangled the subspaces using our distance correlation loss. Panel c compares the distance correlation measure between subspaces of our two model configurations. Panels d and e present t-SNE visualizations of one baseline model and one trained generative model with the disentanglement loss, respectively.
@@
In our second experiment, our objective was to disentangle ethnicity, camera, and patient-identity within three subspaces of a generative model (Fig. 10). Again the baseline model for the generative model already demonstrated relatively good disentanglement performance compared to the encoder model (Fig. 7). However, the application of our disentanglement loss still demonstrated several effects. For instance, the accuracy improvement for camera decoding decreased by 7% for the ethnicity subspace and by 5% for the identity subspace on average (Fig. 10, a versus b). As a direct consequence of minimizing distance correlation between subspaces, this measure consistently decreased for all subspace combinations (Fig. 7c). The t-SNE visualizations not only qualitatively showed the classification and disentanglement performance (Fig. 7d,e), but also pointed us to a correlation in the ethnicity subspace between Indian patients (first row, first column) and the Canon CR-2 AF camera (first row, second column), as they both co-occur in the same clusters, even after incorporating our disentanglement loss. Another interesting finding was that the decoding performance of the ethnicity subspace for ethnicity exhibited a 5% improvement for the generative model compared to the encoder model (Fig. 7a,b). This disparity may suggest that the generative model serves as a better feature extractor because of a better learning bias or a greater model capacity.
@@ 4.4 Realistic image generation and reconstruction
Next, we evaluated the image quality and the reconstruction performance of our generative model. As a baseline we chose a standard StyleGAN2 model with GAN-inversion losses, which we compared to models incorporating our additional losses for independent subspace learning. The baseline model was trained with a 32-dimensional latent space and a single mapping network. For the comparison models, we chose our $\mathcal{W}$-spaces to be either 32- or 36-dimensional for the age-camera-identity and ethnicity-camera-identity setups, respectively. All models generated retinal fundus images with a resolution of $256\times 256$ pixels.
@@
Furthermore, we examined how additional subspace losses affected the image reconstruction performance, by reporting our inversion losses $L_{w}$ and $L_{p}$ for image encoding and pixel space reconstruction, respectively. As expected, image encoding $L_{w}$ consistently showed a slight degradation for models that we jointly optimized with our disentanglement loss (Tab. 1, model C and E). In contrast, the pixel space reconstruction loss $L_{p}$ remained comparable across all configurations. However, $L_{p}$ showed worse mean performance with high variances for some model configurations (Tab. 1, model A and D), suggesting that this loss could potentially benefit from a higher weight, which was also visible in the qualitative reconstruction difference maps in pixel space (Fig. 11d-f).
@@ 4.5 Controllable fundus image generation
In our final experiment, we explored how disentangled subspaces of a generative model influence the generation of retinal fundus images. We started by embedding the original images into latent spaces using our generative model. Because we trained our models with a disentanglement loss, we were able to split the latent spaces of our models into three independent subspaces: the first subspace represented a patient attribute (age or ethnicity), the second subspace encoded the camera, and the third subspace represented patient-identity information. Subsequently, we performed subspace swapping by exchanging the subspace embeddings of different retinal images, allowing us to observe the effect of each subspace on image generation.
@@
Figure 12: Reconstructions of swapped latent subspaces for age-camera-identity disentanglement. In panel a, original fundus images are presented alongside their respective reconstructions in panels b-d. Each column in panels b-d corresponds to the exchange of age, camera, or identity subspaces between the image encodings of panel a. The main diagonals in panels b-d show reconstructions of the images in a, while the off-diagonals illustrate reconstructions where one subspace has been replaced with the subspace indicated in the column header.
For the qualitative analysis, we selected the best of our age-camera-identity models based on disentanglement performance. The original images were randomly sampled from the test set, with the following age and camera labels (Fig. 12, a): (1) 22, Topcon NW 400, (2) 52, Optovue iCam, (3) 76, Canon CR-1. By swapping different subspaces, we generated matrices of image reconstructions from different latent subspace combinations (Fig. 12b-d). Swapping the age subspace between image encodings for reconstruction led to subtle feature changes (Fig. 12b). Notably, bright and reflective features around the thickest vasculature were observed, which tended to fade for age subspaces belonging to an older person (Fig. 12b, first row). These bright and sheen structures are particularly common in the young population and tend to decrease with age (Williams, 1982). We also investigated the impact of changes in the camera subspace on fundus image generation (Fig. 12c). In this model, each column representing a different camera subspace displayed subtle feature changes, occasionally affecting identity features such as the optic disc and vasculature (second column, Fig. 12c). Finally, we swapped the identity subspaces, which were the largest subspaces with 16 dimensions each (Fig. 12d). The idea was to store all remaining information relevant to the image reconstruction in this subspace. As intended, we observed a similar appearance of vasculature, optic disc, and fundus pigmentation in each patient-identity column (Fig. 12d, columns).
@@
Figure 13: Reconstructions of swapped latent subspaces for ethnicity-camera-identity disentanglement. In panel a, original fundus images are presented alongside their respective reconstructions in panels b-d. Each column in panels b-d corresponds to the exchange of ethnicity, camera, or identity subspaces between the image encodings of panel a. The main diagonals in panels b-d show reconstructions of the images in a, while the off-diagonals illustrate reconstructions where one subspace has been replaced with the subspace indicated in the column header.
In a second experiment, we explored the performance of subspace swaps using our best ethnicity-camera-identity disentanglement model (Fig. 13). For this experiment, we randomly sampled three patients from the test set with the following ethnicity and camera labels (Fig. 13a): (1) African, Topcon NW 400, (2) Caucasian, Optovue iCam, (3) Latin American, Canon CR-1. Swapping ethnicity subspaces had a qualitatively stronger visual effect on the image generation than swapping age subspaces (Fig. 13b). In this case, the fundus pigmentation changed for different ethnicity subspaces, which is biologically plausible given the association between retinal pigmentation and ethnicity (Rajesh et al., 2023). However, some ethnicity subspace swaps also altered patient-identity features, such as the appearance of the vasculature (Fig. 13b, last column). The camera subspace swaps exhibited only subtle visual feature changes, occasionally affecting identity features like the vasculature or the optic disc (Fig. 13c, first column). When we changed the identity subspace, identity features (vasculature, optic disc) changed as intended (Fig. 13d, rows).
@@ 5 Discussion
While our model shows excellent image generation performance with disentangled latent representations, some limitations should be acknowledged. One limitation is that the image reconstructions of our generative model showed deficits in capturing fine details, which is particularly evident in the accuracy of the vessel reconstructions. Interestingly, inherent correlations within the dataset pose a challenge for any disentanglement setup. Our experiments revealed that disentangling age from the camera was more straightforward than disentangling ethnicity from the camera, and that performance was limited by the strong correlation between ethnicity and camera in the EyePACS dataset (Träuble et al., 2021; Funke et al., 2022; Roth et al., 2023). In addition, it was challenging to disentangle the free identity space from the camera information. This difficulty may stem from the fact that these represent the largest subspace combinations, spanning 12- and 16-dimensional spaces, thus posing the greatest challenge for distance correlation estimation. Additionally, it is conceivable that camera features are correlated with other dataset features relevant to retinal fundus image generation, further complicating disentanglement. In general, approaches such as ours depend on the dataset distribution at hand and are therefore limited by inherent correlations. For many retrospective data sets obtained from routine care, patient demographics are determined by the visitors to the hospital. With our approach, we offer an algorithmic solution to disentangle (correlated) factors in the dataset to a certain degree.
@@
To some extent, our method is scalable. Our method can be trained on larger datasets because the model training time scales linearly with the size of the dataset. In addition, our method can handle images of different resolutions well, as we show in the Appendix with Fig. F.3. Furthermore, our method can also handle combinations of multiple attributes well, as we show in Fig. G.4. Of course, the combinatorial complexity will eventually prohibit more attributes.
In this study, our focus was exclusively on healthy fundus images due to the additional complexity that would have been introduced by the reconstruction of disease features. However, the issue of spurious correlation between disease and technical features is a practical challenge that we aim to tackle in future work. An additional latent disease subspace, e.g. learned with the EyePACS diabetic retinopathy (DR) labels, could be an interesting clinical test scenario for our work. Here, we could investigate whether a disentangled DR subspace is more robust for disease prediction on a shifted test distribution, e.g. from another hospital. However, in this study we only worked with healthy fundus images due to the unbalanced nature of the dataset. In another study (Ilanchezian et al., 2023), we used diffusion models for counterfactual image generation (without learning any representations), and disease features could be generated here with massive oversampling of the diseased classes, so this may be a way forward here as well. Moreover, alternative measures to distance correlation exist, such as maximum mean discrepancy (MMD) (Sejdinovic et al., 2013), MI bounds, or the use of adversarial classifiers, which can offer advantages, especially in terms of batch size or convergence time. We intend to investigate and compare these methods in the context of subspace disentanglement for medical images in future studies. In this work, we relied on labeled data to encode information into subspaces. Therefore, another compelling avenue for future research is to explore weakly supervised learning approaches, where we could apply a classification loss only on a subset of the training data, in addition to our disentanglement loss. This exploration could potentially yield effective results even on minimally labeled data. Lastly, the main contribution of this work is to extend an inverted GAN in general as an independent subspace learner that is applicable beyond the domain of fundus images. However, since fundus images have very specific properties (circular shape, fine vascular tree structure, optic disc, macula, etc.), biologically plausible biases could improve fundus image generation, which we leave for future work.

### v3

@@ 5.1 Learning disentangled subspaces with an encoder model
Figure 5: Encoder disentanglement performance for age-camera setup. We trained encoder models on retinal images, where the first subspace $w_{\text{age}}$ encoded age $y_{\text{age}}$, and the second subspace $w_{\text{cam}}$ encoded camera information $y_{\text{cam}}$. In panels a and b, we show the accuracy improvement over chance level accuracy. As a baseline model in a, we only trained linear classifiers on top of subspaces. In b, we further disentangled the subspaces using our distance correlation loss. Panel c shows the distance correlation measure between subspaces for both the baseline method and our proposed method for subspace disentanglement. In panels d and e, we provide t-SNE visualizations comparing a baseline model with a model with disentanglement loss, respectively.
Our experiments showed that the age and camera subspaces were heavily entangled in the baseline setting when disentanglement was not enforced (Fig. 5a), such that camera information was present in the age subspace and vice versa. Entanglement was visible by relatively high classification accuracy across subspaces (Fig.5a, off-diagonals). For example, camera type could be decoded from the age subspace $w_{\text{age}}$ with a performance of $28\%$ above chance level, compared to $45\%$ in the camera subspace. When enforcing disentanglement by additionally minimizing the distance correlation between subspaces the cross-subspace classification performance (off-diagonals in Fig. 5b) dropped, while the within-subspace classification was preserved (diagonals). In addition, the distance correlation between subspaces on the test data dropped as a direct consequence of our disentanglement loss (Fig. 5c). We could also observe the disentanglement effect qualitatively in 2D t-SNE visualizations of the learned representations. In both the baseline and the disentangled model, the age subspace exhibited structure w.r.t. age, and the camera subspace encoded camera type well (diagonals in Fig. 5d, e). In the baseline model, however, the age subspace also strongly encoded camera information (first row, second column in d).
Figure 6: Encoder disentanglement performance for ethnicity-camera setup. We trained encoder models on retinal images, where the first subspace $w_{\text{ethn}}$ encoded ethnicity $y_{\text{ethn}}$, and the second subspace $w_{\text{cam}}$ encoded camera information $y_{\text{cam}}$. In panels a and b, we show accuracy improvement over chance level accuracy. As a baseline model in a, we only trained linear classifiers on top of subspaces. In b, we further disentangled the subspaces using our distance correlation loss. Panel c compares the distance correlation measure between subspaces of our two model configurations. In panels d and e, we provide t-SNE visualizations comparing a baseline model with a model with disentanglement loss, respectively.
@@ 5.2 Ablation study for disentanglement loss
In our last ablation experiment, we explored the loss weighting term $\lambda_{DC}$ for distance correlation minimization (Fig. 7c1-c3). As expected, the distance correlation measure between subspaces dropped with larger $\lambda_{DC}$ (Fig. 7c1). At the same time, the decoding accuracy for age from the age subspace (Fig. 7c2, blue curve) collapsed for $\lambda_{DC}\geq 2$, leading to a drop in decoding performance below 10%. In contrast, camera decoding (Fig. 7c2, green curve) from the age subspace becomes less accurate as soon as the disentanglement loss is introduced, indicating that the proposed method does what it is supposed to. For the camera subspace in (Fig. 7c3), the camera decoding accuracy remains high (green), and the age decoding accuracy initially drops (blue), in line with the goals of the distance correlation loss. However, the ablation study showed that the hyperparameter optimization of $\lambda_{DC}$ is essential, as over-weighting can lead to a complete collapse of the subspace encoding (Fig. 7c2,c3).
@@ 5.3 Learning disentangled subspaces with a generative model
As with the predictive model, we encoded label information into latent subspaces by optimizing linear classifiers on distinct tasks. However, in the generative framework, we also introduced an additional latent subspace dedicated to encoding image style, $w_{\text{style}}\in\mathbb{R}^{16}$ (Fig. 4). To ensure independence between all subspaces, we incorporated distance correlation minimization. Consequently, we aimed to minimize the distance correlation between three different subspace combinations. Notably, while we did not have an additional classification head for the style space, we specifically enforced independence from the other subspaces, ensuring that it remained uncorrelated with age, ethnicity, or camera information.
@@
Figure 8: Disentanglement performance for age-camera setup. We trained generative models on retinal images, featuring three subspaces: $w_{\text{age}}$ encoding the age class $y_{\text{age}}$, $w_{\text{cam}}$ encoding the camera class $y_{\text{cam}}$, and $w_{\text{style}}$ serving as a style subspace. In a and b, we report kNN classifier accuracy improvement over chance level accuracy. The baseline model in a involved training linear classifiers on the first two subspaces. In b, we additionally disentangled the subspaces using our distance correlation loss. Panel c compares the distance correlation measure between subspaces of our two model configurations. Panels d and e present t-SNE visualizations of one baseline model and one trained generative model with the disentanglement loss, respectively.
Our disentanglement loss also showed an effect when applied to a generative image model (Fig. 8). While the baseline model for the generative model already demonstrated relatively good disentanglement performance, the introduction of the distance correlation loss further enhanced the separation of subspaces. For example, for the baseline model, age information could be decoded from the camera subspace and vice versa (Fig. 8a, off-diagonals). Additionally, age, and particularly camera information, were still contained within the style subspace of the baseline model (last row, Fig. 8a). Upon incorporating the distance correlation loss, cross-subspace classification performance (off-diagonals and last row, Fig. 8b) dropped, while within-subspace classification was preserved (diagonals). Consequently, the distance correlation between subspaces on the test data decreased (Fig. 8c), highlighting the effect of our disentanglement loss. The qualitative impact of disentanglement was further evident in t-SNE visualizations: for example, the camera labels clustered in the style space for the baseline model, but the clustering became more mixed up when we incorporated the distance correlation loss (Fig. 8d,e).
Figure 9: Disentanglement performance for ethnicity-camera setup. We trained generative models on retinal images, featuring three subspaces: $w_{\text{ethn}}$ encoding the ethnicity class $y_{\text{ethn}}$, $w_{\text{cam}}$ encoding the camera class $y_{\text{cam}}$, and $w_{\text{style}}$ serving as the style subspace. In a and b, we report kNN classifier accuracy improvement over chance level accuracy. The baseline model in a involved training linear classifiers on the first two subspaces. In b, we additionally disentangled the subspaces using our distance correlation loss. Panel c compares the distance correlation measure between subspaces of our two model configurations. Panels d and e present t-SNE visualizations of one baseline model and one trained generative model with the disentanglement loss, respectively.
@@
In our second experiment, our objective was to disentangle ethnicity, camera, and style within three subspaces of a generative model (Fig. 9). Again the baseline model for the generative model already demonstrated relatively good disentanglement performance compared to the encoder model (Fig. 6). However, the application of our disentanglement loss still demonstrated several effects. For instance, the accuracy improvement for camera decoding decreased by 7% for the ethnicity subspace and by 5% for the style subspace on average (Fig. 9, a versus b). As a direct consequence of minimizing distance correlation between subspaces, this measure consistently decreased for all subspace combinations (Fig. 9c). The t-SNE visualizations not only qualitatively showed the classification and disentanglement performance (Fig. 9d,e), but also pointed us to a correlation in the ethnicity subspace between Indian patients (first row, first column) and the Canon CR-2 AF camera (first row, second column), as they both co-occur in the same clusters, even after incorporating our disentanglement loss. Another interesting finding was that the ethnicity decoding performance from the ethnicity subspace exhibited a 5% improvement for the generative model compared to the encoder model (Fig. 6a,b). This disparity may suggest that the generative model serves as a better feature extractor because of a better learning bias or a greater model capacity.
@@ 5.4 Realistic image generation and reconstruction
Next, we evaluated the image quality and the reconstruction performance of our generative model. As a baseline we chose a standard StyleGAN2 model with GAN-inversion losses, which we compared to models incorporating our additional losses for independent subspace learning. The baseline model was trained with a 32-dimensional latent space and a single mapping network. For the comparison models, we chose our $\mathcal{W}$-spaces to be either 32- or 36-dimensional for the age-camera-style and ethnicity-camera-style setups, respectively. All models generated retinal fundus images with a resolution of $256\times 256$ pixels.
@@
Furthermore, we examined how additional subspace losses affected the image reconstruction performance, by reporting our inversion losses $L_{w}$ and $L_{p}$ for image encoding and pixel space reconstruction, respectively. As expected, the image encoding loss $L_{w}$ consistently showed a slight degradation for models that we jointly optimized with our disentanglement loss (Tab. 1, model C and E). In contrast, the pixel space reconstruction loss $L_{p}$ remained comparable across all configurations. However, $L_{p}$ showed worse mean performance with high variances for some model configurations (Tab. 1, model A and D), suggesting that this loss could potentially benefit from a higher weight, which was also visible in the qualitative reconstruction difference maps in pixel space (Fig. 10d-f).
@@ 5.5 Controllable fundus image generation
In our final experiment, we explored how disentangled subspaces of a generative model influence the generation of retinal fundus images. We started by embedding the original images into latent spaces using our generative model. Because we trained our models with a disentanglement loss, we were able to split the latent spaces of our models into three independent subspaces: the first subspace represented a patient attribute (age or ethnicity), the second subspace encoded the camera, and the third subspace represented style information. Subsequently, we performed subspace swapping by exchanging the subspace embeddings of different retinal images, allowing us to observe the effect of each subspace on image generation.
@@
Figure 11: Reconstructions of swapped latent subspaces for age-camera-style disentanglement. In panel a, original fundus images are presented alongside their respective reconstructions in panels b-d. Each column in panels b-d corresponds to the exchange of age, camera, or style subspaces between the image encodings of panel a. The main diagonals in panels b-d show reconstructions of the images in a, while the off-diagonals illustrate reconstructions where one subspace has been replaced with the subspace indicated in the column header.
For the qualitative analysis, we selected the best of our age-camera-style models based on disentanglement performance. The original images were randomly sampled from the test set, with the following age and camera labels (Fig. 11, a): (1) 22, Topcon NW 400, (2) 52, Optovue iCam, (3) 76, Canon CR-1. By swapping different subspaces, we generated matrices of image reconstructions from different latent subspace combinations (Fig. 11b-d). Swapping the age subspace between image encodings for reconstruction led to subtle feature changes (Fig. 11b). Notably, bright and reflective features around the thickest vasculature were observed, which tended to fade for age subspaces belonging to an older person (Fig. 11b, first row). These bright and sheen structures are particularly common in the young population and tend to decrease with age (Williams, 1982). We also investigated the impact of changes in the camera subspace on fundus image generation (Fig. 11c). In this model, each column representing a different camera subspace displayed subtle feature changes, occasionally affecting identity features such as the optic disc and vasculature (second column, Fig. 11c). Finally, we swapped the style subspaces, which were the largest subspaces with 16 dimensions each (Fig. 11d). The idea was to store all remaining information relevant to the image reconstruction in this subspace. As intended, we observed a similar appearance of vasculature, optic disc, and fundus pigmentation in each style column (Fig. 11d, columns).
@@
Figure 12: Reconstructions of swapped latent subspaces for ethnicity-camera-style disentanglement. In panel a, original fundus images are presented alongside their respective reconstructions in panels b-d. Each column in panels b-d corresponds to the exchange of ethnicity, camera, or style subspaces between the image encodings of panel a. The main diagonals in panels b-d show reconstructions of the images in a, while the off-diagonals illustrate reconstructions where one subspace has been replaced with the subspace indicated in the column header.
In a second experiment, we explored the performance of subspace swaps using our best ethnicity-camera-style model (Fig. 12). For this experiment, we randomly sampled three patients from the test set with the following ethnicity and camera labels (Fig. 12a): (1) African, Topcon NW 400, (2) Caucasian, Optovue iCam, (3) Latin American, Canon CR-1. Swapping ethnicity subspaces had a qualitatively stronger visual effect on the image generation than swapping age subspaces (Fig. 12b). In this case, the fundus pigmentation changed for different ethnicity subspaces, which is biologically plausible given the association between retinal pigmentation and ethnicity (Rajesh et al., 2023). Interestingly, some ethnicity subspace swaps also altered other patient features, such as the appearance of the vasculature (Fig. 12b, last column). The camera subspace swaps exhibited only subtle visual feature changes, occasionally affecting patient features like the vasculature or the optic disc (Fig. 12c, first column). When we changed the style subspace, other patient features (vasculature, optic disc) changed as expected (Fig. 12d, rows).
@@ 6 Discussion
While our model shows excellent image generation performance with disentangled latent representations, some limitations should be acknowledged. One limitation is that the image reconstructions of our generative model showed deficits in capturing fine details, which is particularly evident in the accuracy of the vessel reconstructions. Interestingly, we also found that inherent correlations within the dataset pose a challenge for disentanglement. Our experiments revealed that disentangling age from the camera was more straightforward than disentangling ethnicity from the camera, and that the disentanglement performance seemed to be limited by the strong correlation between ethnicity and camera in the EyePACS dataset (Träuble et al., 2021; Funke et al., 2022; Roth et al., 2023). In addition, it was challenging to disentangle the style space from the camera information. This difficulty may stem from the fact that these represent the largest subspace combinations, spanning 12- and 16-dimensional spaces, thus posing the greatest challenge for distance correlation estimation. Additionally, it is conceivable that camera features are correlated with other dataset features relevant to retinal fundus image generation, further complicating disentanglement. In general, approaches such as ours depend on the dataset distribution at hand and are therefore limited by inherent correlations. For many retrospective datasets obtained from routine care, patient demographics are determined by the visitors to the hospital. With our approach, we offer an algorithmic solution to disentangle (correlated) factors in the dataset to a certain degree.
@@
To some extent, our method is scalable. Our method can be trained on larger datasets because the model training time scales linearly with the size of the dataset. In addition, our method can handle images of different resolutions well, as we show in the Appendix with Fig. F.3. Furthermore, our method can also handle combinations of multiple attributes, as we show in Fig. G.4. Of course, the combinatorial complexity will eventually prohibit more attributes.
In this study, our focus was exclusively on healthy fundus images due to the additional complexity that would have been introduced by the reconstruction of disease features. However, the issue of spurious correlation between disease and technical features is a practical challenge that we aim to tackle in future work. An additional latent disease subspace, e.g. learned with the EyePACS diabetic retinopathy (DR) labels, could be an interesting clinical test scenario for our work. Here, we could investigate whether a disentangled DR subspace is more robust for disease prediction on a shifted test distribution, e.g. from another hospital. However, in this study we only worked with healthy fundus images due to the imbalanced dataset. In another study (Ilanchezian et al., 2023), we used diffusion models for counterfactual image generation (without learning any representations), and disease features could be generated here with massive oversampling of the diseased classes, so this may be a way forward here as well.
Moreover, alternative measures to distance correlation exist, such as maximum mean discrepancy (MMD) (Sejdinovic et al., 2013), MI bounds, or the use of adversarial classifiers, which can offer advantages, especially in terms of batch size or convergence time. We intend to investigate and compare these methods in the context of subspace disentanglement for medical images in future studies.
In this work, we relied on labeled data to encode information into subspaces. Therefore, another compelling avenue for future research is to explore weakly supervised learning approaches, where we could apply a classification loss only on a subset of the training data, in addition to our disentanglement loss. This exploration could potentially yield effective results even on minimally labeled data.
Lastly, the main contribution of this work is to extend an inverted GAN in general as an independent subspace learner that is applicable beyond the domain of fundus images. However, since fundus images have very specific properties (circular shape, fine vascular tree structure, optic disc, macula, etc.), biologically plausible biases could improve fundus image generation, which we leave for future work.